# A Quadratic Programming Approach to Flight Envelope Protection Using Control Barrier Functions

### v1

@@ 5.1 Nominal single operating point FEP performance
Figure 9 presents simulation results for the three control cases. The left subplot shows the time series of the angle of attack $\alpha(t)$, where a reference command of $\alpha_{r}(t)=15^{\circ}$ is applied. For illustration purposes, the upper bound on the angle of attack is artificially constrained to $\alpha_{\max}=10^{\circ}$. The green trajectory in Figure 9, representing the aggressive controller, shows significantly faster tracking performance but exhibits overshoot during the transient phase, briefly violating the safety constraint. The blue trajectory representing the closed-loop system with the conservative controller yields a slow reference tracking with no overshoot, suggesting guarantees on FEP when a matching controller setting is found. However, such controller settings will ultimately result in a conservative controller with limited tracking capabilities. To overcome this issue, a safety filter can be used. The orange trajectory illustrates the outcome of applying a CBF-based safety filter that dynamically adjusts the control input from the aggressive controller, as proposed in Eq. (32). The integrated safety filter ensures strict constraint satisfaction, independent of the chosen controller, without compromising performance when the state is comfortably within‚ the safe set. The corresponding phase plot on the right of Fig. 9 further illustrates the described control characteristics of the different controller configurations.
@@ 5.2 Single Operating Point FEP Performance under Disturbances and Parametric Uncertainties
While the previous results focused on the nominal performance of FEP strategies, the following part investigates the robustness of CBF-based safety filters in the presence of external disturbances and model uncertainties. It is important to note that although CBFs are a powerful model-based tool that provides formal guarantees on forward invariance, these guarantees fundamentally depend on the accuracy of the underlying system model. In real-world scenarios, however, discrepancies between the model and the actual system behavior, caused by disturbances, degradation, or unmodeled dynamics, can cause the system to violate safety constraints. This limitation is inherent in all model-based techniques. Adaptive CBF approaches, such as those proposed in [16, 32], aim to restore safety guarantees through online adaptation of model parameters. However, the integration of such adaptive schemes is beyond the scope of this study. Instead, during the analysis of the robustness properties, it was observed that standard CBF safety filters are unable to maintain FEP when the system is subject to external disturbances or significant parametric uncertainties. To address this issue, we utilize the inherent connection between CBFs and Control Lyapunov Functions (CLFs), as proposed in a recent methodology by [33]. Specifically, in situations where the system leaves the safe set, the barrier function $h_{i}(x)$ associated with the respective HOCBF constraint can be reinterpreted as a stabilizing objective by defining an auxiliary CLF:

### v2

@@ 5.1 Nominal single operating point FEP performance
Figure 9 presents simulation results for the three control cases. The left subplot shows the time series of the angle of attack $\alpha(t)$, where a reference command of $\alpha_{r}(t)=15^{\circ}$ is applied. For illustration purposes, the upper bound on the angle of attack is artificially constrained to $\alpha_{\max}=10^{\circ}$. The green trajectory in Figure 9, representing the aggressive controller, shows significantly faster tracking performance but exhibits overshoot during the transient phase, briefly violating the safety constraint. The blue trajectory representing the closed-loop system with the conservative controller yields a slow reference tracking with no overshoot, suggesting guarantees on FEP when a matching controller setting is found. However, such controller settings will ultimately result in a conservative controller with limited tracking capabilities, also in flight envelope regions where no compromise on control performance is needed for safety reasons. To overcome this issue, a safety filter can be used. The orange trajectory illustrates the outcome of applying a CBF-based safety filter that dynamically adjusts the control input from the aggressive controller, as proposed in Eq. (32). The integrated safety filter ensures strict constraint satisfaction, independent of the chosen controller, without compromising performance when the state is comfortably within‚ the safe set. The corresponding phase plot on the right of Fig. 9 further illustrates the described control characteristics of the different controller configurations.
@@ 5.2 Single Operating Point FEP Performance under Disturbances and Parametric Uncertainties
While the previous results focused on the nominal performance of FEP strategies, the following part investigates the robustness of CBF-based safety filters in the presence of external disturbances and model uncertainties. It is important to note that although CBFs are a powerful tool that provides formal guarantees on forward invariance, these guarantees fundamentally depend on the accuracy of the underlying system model. In real-world scenarios, however, discrepancies between the model and the actual system behavior, caused by disturbances, degradation, or unmodeled dynamics, can cause the system to violate safety constraints. This limitation is inherent in all model-based techniques. Adaptive CBF approaches, such as those proposed in [16, 32], aim to restore safety guarantees through online adaptation of model parameters. However, the integration of such adaptive schemes is beyond the scope of this study. Instead, during the analysis of the robustness properties, it was observed that standard CBF safety filters are unable to maintain FEP when the system is subject to external disturbances or significant parametric uncertainties. To address this issue, we utilize the inherent connection between CBFs and Control Lyapunov Functions (CLFs), as proposed by [33]. Specifically, in situations where the system leaves the safe set, the barrier function $h_{i}(x)$ associated with the respective HOCBF constraint can be reinterpreted as a stabilizing objective by defining an auxiliary CLF: